OpenStudy (anonymous):

Please help me! Will give medal :) Which set of data shows forty percent as the experimental probability of a particular traffic light being green on your way to school in the morning?

OpenStudy (anonymous):

A)ten of twenty-five selections at random from a card deck being a black card, with black representing a green light B)2 tails in twenty coin flips, with tails representing a green light C)2 odd-numbers in 8 rolls of a standard die, with odd numbers representing a green light D)spinning twenty vowels out of forty spins of a spinner with 6 equally divided sections labeled with 3 vowels and 3 consonants

OpenStudy (anonymous):

@phi please help

OpenStudy (phi):

can you write the ratios for each choice A,B,C and D ?

OpenStudy (anonymous):

A)10:20 B)2:20 C)2:8 D)Idk

OpenStudy (phi):

re-read choice A and try again

OpenStudy (anonymous):

10:25

OpenStudy (phi):

now change that to a percent. the first step is multiply "top" and "bottom" by 4

OpenStudy (anonymous):

40/100 0.4=40%

OpenStudy (phi):

yes. btw, D is 20:40 (because a spinner with 6 equally divided sections labeled with 3 vowels and 3 consonants means the chance of a vowel is the same as a chance for a consonant) (in other words, this is very similar to 20 tails out of 40 flips problem)

OpenStudy (anonymous):

ok..thank you so much!! so the answer is A

OpenStudy (phi):

if you work out the other choices, they will not have 40% A)10:25 = 40% B)2:20 = 1/10 = 10% C)2:8 = ¼ = 25% D) 20:40= ½ = 50%

OpenStudy (anonymous):

ohhh,ok.. thanks agian

OpenStudy (phi):

the only way to get good at this stuff, is work at it. I would try to change each choice to a decimal and then a percent and check that I got it right.

OpenStudy (anonymous):

k:)